OpenStudy (anonymous):

Let A = {circle, square} and let B = {3, 4, 5}. Find A X B

OpenStudy (anonymous):

\[A \times B=\left\{ \left( circle,3 \right),\left( circle,4 \right),\left( circle,5 \right), \left(\square,3 \right),\left( \square,4 \right),\left( \square,5 \right) \right\}\]